33. Consider a binomial experiment with n 5 20 and p 5 .70.

PLEASE SHOW ANSWERS AND FORMULAS IN EXCEL

a. Compute f(12).

b. Compute f(16).

c. Compute P(x $16).

d. Compute P(x #15).

e. Compute E(x).

f. Compute Var(x)
Answer #1

Solution:

n=20, p=0.70

a.)

f(12)=0.1144

(use excel function: =BINOMDIST(12,20,0.7,FALSE))

b.)

f(16)=0.1304

(use excel function: =BINOMDIST(16,20,0.7,FALSE))

c.)

d.)

e.)

E(x)=np=20*0.7=14

f.)

Var(x)=np(1-p)=20*0.7*(1-0.7)=4.2